Compute the payback period for each of these two separate investments:

  1. A new operating system for an existing machine is expected to cost $270,000 and have a useful life of five years. The system yields an incremental after-tax income of $77,884 each year after deducting its straight-line depreciation. The predicted salvage value of the system is $10,000.
  2. A machine costs $210,000, has a $14,000 salvage value, is expected to last ten years, and will generate an after-tax income of $43,000 per year after straight-line depreciation.
Answer #1

a) Payback period = Initial investment/Annual cash flow

Annual cash flow = 77884+52000 = 129884

Payback period = 270000/129884 = 2.08 Years

b) Annual cash flow = 43000+19600 = 62600

Payback period = 210000/62600 = 3.35 years